OpenStudy (anonymous):

assume that a random sample is used to estimate a population proportion p. Find the margin of error E that corresponds to the given statistics and confidence level. n=500, x=300, 95 percent confidence

OpenStudy (amistre64):

nice, but not relevant to the question

OpenStudy (anonymous):

thats what i am saying.

OpenStudy (amistre64):

\[E=z\sqrt{\frac{pq}{n}}\]

OpenStudy (amistre64):

another way to look at this is: \[E=z\sqrt{\frac{x(n-x)}{n^3}}\]

OpenStudy (amistre64):

the z score of 95% is what ... 1.96 ?

OpenStudy (amistre64):

does this look familiar?

OpenStudy (anonymous):

yes but i am still confused on how to find p hat. I know q hat is 1-phat

OpenStudy (amistre64):

phat = x/n

OpenStudy (amistre64):

\[\Huge \frac{pq}{n}=\frac{\frac{x}{n}\frac{n-x}{n}}{n}=\frac{x(n-x)}{n^3}\]

OpenStudy (amistre64):

1.96sqrt(300*200/500^3) = .0429414...

OpenStudy (amistre64):

if your not given decimal percents, then to avoid excess calculations I would just use the n^3 version

OpenStudy (anonymous):

thank you i worked it out and got the same

OpenStudy (amistre64):

youre welcome